OpenStudy (anonymous):

7x+15y=32 x-3y=20 solve using elimination

OpenStudy (gorv):

from x-3y=20 x=20+3y put in otther equation 7*(20+3y)+15y=32 140+21y+15y=32 36y=32-140=-102 y=-102/36=-17/6

OpenStudy (gorv):

x=23/2

OpenStudy (anonymous):

thank you gorv

OpenStudy (gorv):

wlcm :)